Example 1: A cantilever beam 3 m-long with a channel section carries a concantrted load P=12 kN at the free hand. The load makes an angle 0-30° with the vertical 200 y-axis and passes through the shear center. Calculate the stresses at A, B, C and D (in MPa). 120 20 12 kN SC 12 kN 3 m
Answer #1

20mm 18mm 2n Uiomm 70mm 130%20 xtra- 82mm = 200-82= {{8mm fLom top manha [20d1903 +180x20 x (118-90 Tx.x 2 +140x20 = 10x 20x72 39.691 m ox20st 180x20x60)x2+ 20 1 2. 千Ay 1xlo 30.33xlo -0,785米11g+ , 586×70 51Gl M Shem at B MM.wje_ My...。6 = -0785x118 + . 586 x70 133. 65mfa o·า 85,82-0586 x70